Evaluate the following indefinite integrals:

  1. ∫ (12x^5 βˆ’ 6π‘₯^3 βˆ’ 4x + 1/2 ) dx
  2. ∫ (√π‘₯ + √3)^2 dx
  3. ∫ (π‘₯^𝑒 βˆ’ π‘₯/𝑒 + 2𝑒π‘₯) dx
  4. ∫ (2𝑑+1)(2π‘‘βˆ’1)/2βˆšπ‘‘ dx
  5. ∫ 8𝑦^3 βˆ’1/2𝑦+1 dx
1
Expert's answer
2021-02-28T17:04:57-0500

Ques:1:

"\intop" "(12x^5-6x^3-4x+\dfrac{1}{2})dx"

Distribute the integeration on each term

"\Rightarrow\smallint12x^5dx-\smallint6X^3dx-\smallint4xdx+\smallint\dfrac{1}{2}dx"

"\Rightarrow12\times\dfrac{x^{(5+1)}}{5+1}-6\times\dfrac{x^{(3+1)}}{3+1}-4\times\dfrac{x^{1+1}}{1+1}+\dfrac{1}{2}\times\dfrac{x^{0+1}}{0+1}"


"\Rightarrow" "12\times\dfrac{x^6}{6}-6\times\dfrac{x^4}{4}-4\times\dfrac{x^2}{2}+\dfrac{1}{2}\times x"


"\Rightarrow2x^6-\dfrac{3}{2}x^4-2x^2+\dfrac{x}{2}"



Ques :2:

"\smallint(\sqrt{x}+\sqrt{3})^2dx"


"\Rightarrow\smallint( x+3+2\sqrt{3}\sqrt{x})dx"


"\Rightarrow \smallint xdx+3\smallint dx+\smallint2\sqrt{3}\times x^\dfrac{1}{2}dx"


"\Rightarrow \dfrac{x^2}{2}+3x+2\sqrt{3}\times\dfrac{x^{\dfrac{1}{2}+1}}{\dfrac{1}{2}+1}"


"\Rightarrow\dfrac{x^2}{2}+3x+2\sqrt{3}\times \dfrac{x^{\dfrac{3}{2}}}{\dfrac{3}{2}}"


Ques::3::


"\smallint( x^e-\dfrac{x}{e}+2\times e \times x)dx"


"\Rightarrow\smallint x^edx-\smallint \dfrac{x}{e}dx+\smallint2exdx"


"\Rightarrow \dfrac{x^{e+1}}{e+1}-\dfrac{x^2}{2e}+ex^2"


Ques::4::


"\smallint\dfrac{(2t+1)(2t-1)}{2\sqrt{t}}dx"


As we have to integerate in terms of "x" .so the terms containing "t" can be taken outside the integeration


"\Rightarrow" "\dfrac{(2t+1)(2t-1)}{2\sqrt{t}}\smallint dx"


"\Rightarrow\dfrac{(4t^2-1)}{2\sqrt{t}}\times x"


Ques::5::


"\smallint(8y^3-\dfrac{1}{2y}+1)dx"


"\Rightarrow8y^3\smallint dx- \dfrac{1}{2y}\smallint dx+\smallint dx"


"\Rightarrow 8y^3\times x-\dfrac{x}{2y}+x"